Required information

[The following information applies to the questions displayed below.]

Vernon Company is a retail company that specializes in selling outdoor camping equipment. The company is considering opening a new store on October 1, 2019. The company president formed a planning committee to prepare a master budget for the first three months of operation. As budget coordinator, you have been assigned the following tasks:

Required

  1. October sales are estimated to be $280,000, of which 45 percent will be cash and 55 percent will be credit. The company expects sales to increase at the rate of 25 percent per month. Prepare a sales budget.

  2. The company expects to collect 100 percent of the accounts receivable generated by credit sales in the month following the sale. Prepare a schedule of cash receipts.

  3. The cost of goods sold is 60 percent of sales. The company desires to maintain a minimum ending inventory equal to 10 percent of the next month’s cost of goods sold. However, ending inventory of December is expected to be $13,600. Assume that all purchases are made on account. Prepare an inventory purchases budget.

  4. The company pays 80 percent of accounts payable in the month of purchase and the remaining 20 percent in the following month. Prepare a cash payments budget for inventory purchases.

  5. Budgeted selling and administrative expenses per month follow:

Salary expense (fixed) $ 19,600
Sales commissions 5 % of Sales
Supplies expense 2 % of Sales
Utilities (fixed) $ 3,000
Depreciation on store fixtures (fixed)* $ 5,600
Rent (fixed) $ 6,400
Miscellaneous (fixed) $ 2,800
  1. *The capital expenditures budget indicates that Vernon will spend $237,600 on October 1 for store fixtures, which are expected to have a $36,000 salvage value and a three-year (36-month) useful life.

Use this information to prepare a selling and administrative expenses budget.

  1. Utilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in which they are incurred. Prepare a cash payments budget for selling and administrative expenses.

  2. Vernon borrows funds, in increments of $1,000, and repays them on the last day of the month. Repayments may be made in any amount available. The company also pays its vendors on the last day of the month. It pays interest of 1 percent per month in cash on the last day of the month. To be prudent, the company desires to maintain a $28,000 cash cushion. Prepare a cash budget.
Answer #1

sales budget Estimated Total Sales Estimated Cash Sales Estimated Credit Sales October November December $ 280,000 $ 350,000DE F cash payments budget for inventory purchases October November December For current month purchases $ 151,200 $ 172,200 $B49 D E F G cash budget October November December $0 $ 28,800 $ 28,000 $ 126,000 $311,500 $ 389,375 $ 126,000 $340,300 $ 417,

for formulas and calculations, refer to the image below -

A F a s Estimated Total Sales Estimated Cash Sales Estimated Credit Sales ales budget October 280000 =D4*45% =D4*55% NovemberB22 А в с D Ε F cash payments budget for inventory purchases October November For current month purchases =D20*80% =E20*80% FB49 00 60 cash budget October November =D63 =E12 =E51+E52 =D12 December =E63 =F12 =F51+F52 =D51+052 beginning balance add: ca

In case you have any query, kindly ask in comments.